A weight A lbs. on one side of a beam balances a weight of 40 lbs. placed 6ft from the fulcrum on the other side. If the unknown weight is moved 3ft nearer the fulcrum, it balances a weight of 20lbs placed ft from 7 1/2 the fulcrum. Find the unknown weight and its distance from the fulcrum in the first instance. (Neglect the weight of the beam.)


1
Expert's answer
2022-04-05T15:17:36-0400

Define:

x lbs - unknown weight

y ft - distance from the fulcrum in the first instance

We can make the following system of equations:

"x\cdot y=40\cdot 6"

"x(y-3)=20\cdot 7\frac12"

Let’s solve it. From the second equation:

"xy-3x=20\cdot 7\frac12=150"

Putting the first equation into the second we will obtain:

"40\cdot6-3x=150"

"3x=90"

"x=30"

"30\cdot y=40\cdot6"

"y=8"

Answer: A weight is 30lbs, a distance is 8ft